Apple and Samsung both make mobile phones and tablets using labour as an input. The amount of labour hours needed to produce one unit of each type of device is as follows: Hours per unit Mobile Phone Tablet 2 5 Apple Samsung 5 7 Let 14 and 14 denote the labour employment by Apple in the production of mobile and tablet, respectively. Similarly, lm and 1 denote the labour employment by Samsung in the production of mobile and tablet, respectively. It is further given that lm + 14 = 70 and le+ 1 = 70. i. Calculate opportunity cost of each type of device for each company and show that Apple has comparative advantage in producing mobile phones. ii. If each company divides equally the labour available in the production of the two devices, what's the total production of each device in the economy? (2) iii. Suggest an allocation of labour for both the companies such that the total production of each device is higher than that in part (ii). (2)


1
Expert's answer
2022-03-30T14:13:33-0400

The opportunity cost for each tablet is

For Apple

1 tablet "= \frac{5}{2}= 2.5" phones


For Samsung

1 tablet "=\frac{7}{5}= 1.4" phones


It is cheaper for Samsung to produce tablets which implies that Apple has a comparative advantage in mobile production.


ii) At Apple

Mobile "= \frac{70}{2\times2} = 17.5"

Tablet "= \frac{70}{2\times5} =7"


At Samsung

Mobile "=\frac{70}{2\times5}=7"

Tablet "= \frac{70}{2\times7} = 5"


Total mobiles "= 24.5"

Total tablet "= 12"